Your teacher is giving you a test worth 100 points and containing 40 questions. There are 2-point and 4-point questions on the test. How many of each type of questions are on the test?
Answer by Earlsdon(6294)

Let x = the number of 2-point questions, then (40-x) = the number of 4-point questions. You can write the equation to solve this:
2(x) + 4(40-x) = 100 Simplify and solve for x.
2x + 160 - 4x = 100
-2x + 160 = 100
-2x = -60
x = 30 and (40-x) = 10
There are thirty 2-point questions and ten 4-point questions.